OpenStudy (anonymous):

x^2+y^2=5 x^2-y^2=2 find all four solutions

OpenStudy (anonymous):

x^2+y^2=5=x^2+y^2-5=0

OpenStudy (anonymous):

x^2-y^2=2=x^2-y^2-2=0

OpenStudy (anonymous):

add these two equations you get:  notice this gives you two x values but you will also get two y- values for a total of 4 points where these two graphs intersect.